A rectangular field has a perimeter of 220 meters and an area of 2,856 square meters. What are its dimensions?

Answer by mananth(16946)

L+W=110
L*W= 2856
..
w=110-L
L*(110-L)=2856
110L-L^2=2856
L^2-110L+2856=0
L^2_68L-42L+2856=0
L(L-68)-42(L-68)=0
(L-68)(L-42)=0
L=62 OR 48
The dimensions are 62 by 48 meters